OpenStudy (anonymous):

How to calculate the limit of (n^(1/n)-1)* ln n as n goes to infinity

terenzreignz (terenzreignz):

\[\lim _{n \rightarrow \infty}[n^{\frac{1}{n} - 1}\ln n]\] Is this what you meant?

OpenStudy (ash2326):

\[\lim _{n \rightarrow 1}[n^{\frac{1}{n} - 1}\ln n]\] I think n\(\to 1\)

OpenStudy (anonymous):

No \[\lim_{n \rightarrow \infty} (\sqrt[n]{n} -1) \ln n\]

terenzreignz (terenzreignz):

Well, as n goes to infinity, ln n goes to infinity and so does the nth root of n, so..

terenzreignz (terenzreignz):

no wait

OpenStudy (anonymous):

The answer should be 0, but I don't know how to get there

terenzreignz (terenzreignz):

Ok, I can get there, but it require knowledge of L'Hopital's rule, are you familiar with it?

OpenStudy (amistre64):

1/n approaches zero n^(1/n) approaches 1 1-1=0 0 lnn = 0 iis my view of it

OpenStudy (amistre64):

Lhop just seems like id be messy :)

terenzreignz (terenzreignz):

Lhop is required to prove the n^(1/n) thing, though. It's indeterminate

OpenStudy (anonymous):

Yeah, definitely, I know L'hop.

terenzreignz (terenzreignz):

Well, as amistre64 said, the limit of n^(1/n) is 1, but can you show it? :)

OpenStudy (anonymous):

Yes, I can do that

terenzreignz (terenzreignz):

so, from there, you'll get (1-1)ln n which is just 0 ln n

OpenStudy (anonymous):

Yeah, this is where I'm stuck. Getting out of the zero multiplied by inf part

terenzreignz (terenzreignz):

It's not 0 times infinity, it's 0 times ln n, which is just 0 :)

OpenStudy (anonymous):

but isn't ln n as n goes to infinity, infinity

terenzreignz (terenzreignz):

the limit of n as n goes to infinity is also infinity But I don't think you're about to say that lim 0n = well, something indeterminate :)

OpenStudy (anonymous):

So, would a sequence like this diverge or converge then?!

terenzreignz (terenzreignz):

oh it's a sequence? I thought it was a function >.> the sequence converges, because the limit exists, in fact, it converges to 0 :)

OpenStudy (anonymous):

Yeah, it's a sequence. Sorry, I guess I should've mentioned that part. Sorry, but I still don't get the part about 0 lnn. So, wnevever I get to 0 multiplied by another part that goes to inifinity, I should just treat inf like any other number and so the multiplication would be zero?

terenzreignz (terenzreignz):

as long as it's plainly 0 and not some function that just goes to 0 as n approaches infinity

OpenStudy (anonymous):